Using the following information for the region of New England:

           Nominal GDP                  Real GDP                            Population                          Year

           (millions)                         (millions 2005 $’s)                          

                   491,139                           623,792                               13,642,253                          1997     

                   799,219                           830,455                               14,279,291                          2007

                 1,021,856                        883,228                               14,810,068                          2017

  1. What is the average annual growth rate for New England’s economy for the time periods listed?
  2. What is the average annual inflation rate for New England for the time periods listed?
  3. To what extent are living standards likely changing in this economy? Support your claim using evidence from the above table. Compare the time periods and comment on any changes.
Answer #1

a. average annual growth rate for New England’s economy for the time periods listed

For a period 1997-2007:

For growth real GDP Values should be taken.

Rate of growth is calculated by a formula: (Final value- Initial value/Initial value)*100

Therefore, (830,455- 623,792/623,792)*100 = (206,633/623,792)*100 = 33.12% For full ten year period and hence average per year growth is : 33.12/10= 3.31%

Similarly for years 2007-2017:

(883,228-830,455/830455)*100= (52773/830455)*100 = 6.35 % for 10 year,

On an average per year growth 0.63%

b. Considering 1997 as base year.

Deflator : (Nominal GDP/ Real GDP)*100 =

for 1997-2007:

(799219/830455)*100 = 96.23

For 2007-2017:

(1021856/883228)*100 = 115.69

Inflation rate between 2007-2017: (Deflator 2017- Deflator 2007/ Deflator 2007)*100:

(115.69-96.23/.96.23)*100 = 20.22 % For 10 years. Therefore per year: 2.02 %

For 1997-2007 it was:

(96.23-100/100)*100= 3.77 % Per year it was: 0.37 %

c. Per year GDP Growth rate average in between: 1997-2007 was

3.31%

In between 2007-2017 it was 0.63%. it means GDP progressed slowly in second decade. Inflation rate was very low in first period as compared to second year. Hence in second period even if nominal GDP shows growth real GDP growth is very very low as price increase has nullified effect of growth.